# AgileFormer: Spatially Agile Transformer UNet for Medical Image Segmentation

### v1

@@ Abstract
In the past decades, deep neural networks, particularly convolutional neural networks, have achieved state-of-the-art performance in a variety of medical image segmentation tasks. Recently, the introduction of the vision transformer (ViT) has significantly altered the landscape of deep segmentation models. There has been a growing focus on ViTs, driven by their excellent performance and scalability. However, we argue that the current design of the vision transformer-based UNet (ViT-UNet) segmentation models may not effectively handle the heterogeneous appearance (e.g., varying shapes and sizes) of objects of interest in medical image segmentation tasks. To tackle this challenge, we present a structured approach to introduce spatially dynamic components to the ViT-UNet. This adaptation enables the model to effectively capture features of target objects with diverse appearances. This is achieved by three main components: (i) deformable patch embedding; (ii) spatially dynamic multi-head attention; (iii) deformable positional encoding. These components were integrated into a novel architecture, termed AgileFormer. AgileFormer is a spatially agile ViT-UNet designed for medical image segmentation. Experiments in three segmentation tasks using publicly available datasets demonstrated the effectiveness of the proposed method. The code is available at https://github.com/sotiraslab/AgileFormer.
Keywords Transformer $\cdot$ UNet $\cdot$ Medical Image Segmentation.
Figure 1: (a) Comparison of model scalability in the Synapse dataset. The base model is almost four times larger than the tiny model. (b) A roadmap going from SwinUNet to the proposed AgileFormer: from the top to the bottom, each row represents a model design variant. The foreground bars are DSC in the FLOP regime of different design variants; a hatched bar means the modification leads to a performance drop.
@@ 1 Introduction
Medical image segmentation tasks are important in modern medicine, as it is typically the first step in many image-driven diagnoses and analyses [1, 2]. Deep learning based automated segmentation methods have dominated this field due to their high efficiency and state-of-the-art performance. Among them, convolutional neural networks (CNNs) [3, 4, 5, 6, 7, 8] have emerged as the most prevalent choice since the introduction of UNet [3]. This is due to CNNs’ inherent advantages in handling image-driven tasks, such as their ability to capture locality and translation invariance. However, they struggle with capturing global semantics primarily due to their restricted receptive field.
@@
In contrast, the recently proposed vision transformer (ViT) [9] mitigates this problem through a self-attention mechanism that captures dependencies between image patches regardless of their spatial distances. The first ViT-based UNet (ViT-UNet) for medical image segmentation combined a ViT encoder with a CNN decoder, and hence was termed TransUNet [1]. However, TransUNet is buderened by a large number of parameters ($\sim$100M) and considerable computational complexity. This is because it employs standard self-attention, which has a quadratic time and memory complexity w.r.t. the input token size. Leveraging the window attention [10], which performs self-attention within a small window in parallel for all image patches, SwinUNet [2] mitigates this challenge and is the first pure ViT-UNet, with self-attention as the main feature extractor. However, SwinUNet uses fixed-size windowing, which may limit its ability to capture precise representation for target objects with varying sizes and shapes. This potentially limits its adaptability and generalizability to multi-class segmentation tasks.
Additionally, SwinUNet do not exhibit good scaling behavior. In other words, their performance did not scale effectively when increasing the size of the model [2] (as evident in Figure˜1(a)). This is in contrast to evidence provided by empirical studies using natural images that have revealed that ViTs exhibit exceptional scaling behavior. We hypothesize this phenomenon may also be attributed to heterogeneity in the size and shape of objects targeted for segmentation, which cannot be handled by fixed-size windowing. To address these challenges, spatially dynamic components are necessary for improving the performance and scalability of a ViT-UNet. Although recent works have explored introducing dynamic (e.g., multi-scale [11], deformable [12, 13]) components into ViT-UNets for medical imaging, they have mainly focused on adopting existing deformable attention. However, they did not investigate introducing dynamic components into other parts (i.e., patch embedding and positional encoding) of a ViT-UNet. We argue that enhancing these other components with dynamic elements is equally important. We kindly direct the readers to Appendix˜A for an in-depth discussion on how AgileFormer fundamentally differs from the three aforementioned works  [11, 12, 13].
@@
Accordingly, we developed a novel spatially agile pure ViT-UNet to capture diverse target objects in a medical image segmentation task. First, we replace the standard rigid square patch embedding in ViT-UNet with a novel deformable patch embedding. Second, we adopt spatially dynamic self-attention [14] as the building block to capture spatially varying features. Third, we propose a novel multi-scale deformable positional encoding to model irregularly sampled grids in self-attention. We integrated these dynamic components into a novel ViT-UNet architecture, named AgileFormer. Extensive experiments demonstrated the effectiveness of the proposed method for three medical image segmentation tasks. AgileFormer outperformed recent state-of-the-art UNet models for medical image segmentation and demonstrated exceptional scalability.
@@ 2 Method
In this section, we provide a roadmap going from the standard SwinUNet to the proposed AgileFormer (see Figure˜1(b)). To make this paper self-contained, we first provide a brief introduction of the essential elements in ViT-UNets. A ViT-UNet (e.g., SwinUNet) is a U-shaped encoder-decoder network with skip connections, wherein the primary feature extractions in both the encoder and decoder are achieved by the self-attention mechanism. The building block of a standard ViT-UNet is composed of three fundamental components: patch embedding, self-attention, and positional encoding. The patch embedding projects image patches into feature embeddings. More recent ViTs [15, 14] even treat downsampling (i.e., patch merging in a SwinUNet) as part of the patch embedding. In this paper, we will follow the same convention. The self-attention, which captures the dependencies between image patches, is used for the main feature extraction. For computational tractability and the need of locality, recent ViT-UNets use window-based self-attention mechanism [2, 16, 13, 17]. Unlike convolution, self-attention discards the spatial correlation between image patches, which hinders the localization ability of pure ViT in segmentation task. The positional encoding is used to address this limitation. We would like to point out that there are also hybrid models [1, 16, 18, 17, 12, 19], where the main feature extraction in the encoder/decoder is achieved by both CNN and ViT. Whereas, we focus on developing a pure ViT-UNet in this paper.
Figure 2: The architecture of the proposed AgileFormer.
The rest of this section is organized to introduce the three main components of the proposed AgileFormer in Figure˜2: deformable patch embedding (Section˜2.1), spatially dynamic multi-head attention (Section˜2.2), and deformable positional encoding (Section˜2.3).
@@ 2.1 Deformable Patch Embedding
2.1.1 Rigid patch embedding
The ViT-UNet starts by converting image patches into tokens. This process typically involves partitioning an image into a sequence of non-overlapping $n\times n\ (\times n)$ (e.g., $4\times 4\ (\times 4)$ in SwinUNet [2, 20, 17]) fixed-size patches. Subsequently, each of these patches is projected into a 1D feature vector. The main reason for performing a rigid (square) patch embedding is due to its simplicity, as it can be easily implemented as a standard convolution (kernel size $k=n$; stride $s=n$; dilation $d=1$). However, we argue that this rigid patch embedding is not optimal for a segmentation task for two main reasons. First, segmentation requires more precise pixel-level localization. However, the rigid patch embedding can only provide patch-level localization. Second, the shape and size of target objects in most medical image segmentation tasks (e.g., the multi-organ segmentation task) vary significantly.
2.1.2 Deformable patch embedding
To address the limitation of rigid patch embedding, we propose a deformable patch embedding (see Figure˜2) by leveraging the deformable convolution [21]. The deformable convolution is defined as
@@ 2.1.2 Deformable patch embedding
where $\boldsymbol{f}\in\mathbb{R}^{L\times d_{f}}$ is a $d_{f}$-dimensional feature map with a uniform grid of $L$ locations $p\in\mathbb{R}^{L\times D}$ ($D=2$ for 2D; $D=3$ for 3D). $\boldsymbol{k}$ is the convolutional kernel that operates on the grid $\Omega=[p_{k}]$ defines $k$-nearest (i.e., kernel size) neighboring locations of $p$. $\Delta p]$ are offsets from which the irregular grid is sampled. The offsets are learned through a single convolutional layer $\Delta p=\operatorname{Conv}_{\text{offset}}(\boldsymbol{f})$. $\phi$ is a sampling function that performs bilinear/trilinear interpolation to sample locations $p+\Delta p+p_{k}]$ in $\boldsymbol{f}$, as the offsets $\Delta p$ are typically fractional.
@@
Downsampling patch embedding. We also replaced the patch merging in standard SwinUNet with convolutional downsampling via a single layer ($k=3$; $s=2$; $d=1$) for downsampling. We used overlapping kernels to better preserve localized patterns [22, 15, 20], which also aligns with the overlapping deformable patch embedding.
@@ 2.2 Spatially Dynamic Self-Attention
The self-attention is the building block of a ViT UNet. Unlike convolution, self-attention does not enforce any spatial inductive bias but makes decisions purely by relying on dependencies (mainly similarities) between tokens while lacking the ability to capture spatially adaptive features for multi-class segmentation. Accordingly, we propose using a spatially dynamic self-attention module as the building block of ViT-UNet. This module is inspired from [15], and it includes deformable multi-head attention (DMSA) [15] and neighborhood multi-head attention (NMSA) [23]. The transformer block is constructed by alternating these two attention mechanisms (see Figure˜2). We also distribute more computation to the third encoder block with a stage ratio of $[1:2:5:1]$, instead of $[2:2:2:2]$ in SwinUNet. This is because the third layer of the encoder typically captures better feature representations than the others [24].
2.2.1 Deformable Multi-head Self-Attention (DMSA)
@@
$\texttt{DMSA}_{h}(\boldsymbol{f})=\operatorname{softmax}(\boldsymbol{Q}_{h}\tilde{\boldsymbol{K}}_{h}^{\top}/\sqrt{d_{k}})\tilde{\boldsymbol{V}}_{h}$ (2)
@@
$\boldsymbol{Q_{h}}=\boldsymbol{f}\boldsymbol{W}_{h}^{Q},\quad\tilde{\boldsymbol{K}}_{h}=\tilde{\boldsymbol{f}}\boldsymbol{W}_{h}^{K},\quad\tilde{\boldsymbol{V}}_{h}=\tilde{\boldsymbol{f}}\boldsymbol{W}_{h}^{V},\quad\tilde{\boldsymbol{f}}=\phi(\boldsymbol{f};p+\Delta p_{h}).$ (3)
@@
Here, we reuse the notation in Equation˜1 with $p$ being a uniform grid of points, $\Delta p_{h}$ being the generated offsets for the $h$-th head, and $\phi$ being an interpolation function. $\{\boldsymbol{W}_{h}^{Q},\boldsymbol{W}_{h}^{K}\}\in\mathbb{R}^{d_{f}\times d_{k}}$ and $\boldsymbol{W}_{h}^{V}\in\mathbb{R}^{d_{f}\times d_{v}}$ are trainable parameters, and $d_{k}$, $d_{v}$ are the hidden dimensions of linear projection of the key $\tilde{\boldsymbol{K}}_{h}$ and the value $\tilde{\boldsymbol{V}}_{h}$ in DMSA. The offsets in DMSA are also generated by passing the query through a convolutional layer $\Delta p=\operatorname{Conv}_{\text{offset}}(\boldsymbol{Q}_{h})$. The resulting irregularly sampled feature map is denoted as $\tilde{\boldsymbol{f}}$ in Equation˜3. Similar to deformable convolution, the irregularly sampled feature map is then applied to self-attention by using irregularly sampled key $\tilde{\boldsymbol{K}}_{h}$ and value $\tilde{\boldsymbol{V}}_{h}$ (see Equation˜2).
2.2.2 Neighborhood Multi-head Self-Attention (NMSA)

### v2

@@ Abstract
In the past decades, deep neural networks, particularly convolutional neural networks, have achieved state-of-the-art performance in a variety of medical image segmentation tasks. Recently, the introduction of vision transformers (ViTs) has significantly altered the landscape of deep segmentation models. The excellent performance and scalability of ViTs have driven a growing focus on their application. However, we argue that the current design of the vision transformer-based UNet (ViT-UNet) segmentation models may not effectively handle the heterogeneous appearance (e.g., varying shapes and sizes) of target objects. In addition, prevailing ViT-UNets focus on enhancing the self-attention building block while neglecting other important components (i.e., patching embedding and positional encoding). To tackle these limitations, we present a structured approach to introduce spatially dynamic components into a ViT-UNet. This enables the model to effectively capture features of target objects with diverse appearances. This is achieved by three main components: (i) deformable patch embedding; (ii) spatially dynamic multi-head attention; (iii) multi-scale deformable positional encoding. These components are integrated into a novel architecture, termed AgileFormer. Experiments in three segmentation tasks using publicly available datasets (Synapse multi-organ, ACDC cardiac, and Decathlon brain tumor datasets) demonstrated the effectiveness of AgileFormer for 2D and 3D segmentation tasks. Remarkably, AgileFormer sets a new state-of-the-art performance with a Dice Score of 85.74% and 87.43 % for 2D and 3D multi-organ segmentation on Synapse without significant computational overhead.
{strip}
Figure 1: Left: Qualitative comparison between the proposed AgileFormer and nnFormer [47] on the Synapse multi-organ segmentation task, where each white dashed box marks inaccurate segmented regions. For demonstration purposes, we visualize all 13 organs, but we only report results for 8 of them. We can observe that nnFormer struggled to accurately segment the spleen and stomach, confusing one for the other. Moreover, it over-segmented the right kidney. This is likely due to the fact that fixed-sizing window attention and patch embedding cannot accurately capture objects with varying sizes and shapes, and hence produce inaccurate feature representations. In contrast, AgileFormer, which can capture spatially varying representations via deformable patch embedding, spatially dynamic self-attention and multi-scale deformable positional encoding, accurately segmented organs with varying sizes and shapes. Right: Segmentation accuracy (DSC) against model complexity (number of parameters and FLOPs) on the Synapse multi-organ segmentation task. In both 2D and 3D settings, AgileFormer outperformed recent state-of-the-art methods, while having fewer parameters and FLOPs.
@@ 1 Introduction
Medical image segmentation tasks are important in modern medicine, as they typically serve as the first step in many image-driven diagnoses and analyses [5, 4]. Deep learning based automated segmentation methods have dominated this field due to their high efficiency and state-of-the-art performance. Among them, convolutional neural networks (CNNs) [30, 26, 20, 18, 17] have emerged as the prevalent choice since the introduction of UNet [30]. This is due to CNNs’ inherent advantages in handling image-driven tasks, such as their ability to capture locality and translation invariance. However, they struggle with capturing global semantics primarily due to their restricted receptive field.
@@
In contrast, the recently proposed vision transformer (ViT) [8] mitigates this problem through a self-attention mechanism that captures dependencies between image patches regardless of their spatial distances. The first ViT-based UNet (ViT-UNet) for medical image segmentation combined a ViT encoder with a CNN decoder, and hence was termed TransUNet [5]. However, TransUNet is buderened by a large number of parameters ($\sim$100M) and considerable computational complexity. This is because it employs standard self-attention, which has quadratic time and memory complexity w.r.t. the input token size. SwinUNet [23] addressed this challenge by leveraging window attention [23], which performs self-attention within small windows in parallel for all image patches. SwinUNet is the first pure ViT-UNet, with self-attention as the main feature extractor for the encoder, decoder, and bottleneck of a UNet. However, SwinUNet uses fixed-size windowing, which may limit its ability to capture precise representations for target objects with varying sizes and shapes. Additionally, previous ViT-UNets deploy a fixed-size patch partitioning (i.e., patch embedding), which inherently limits their ability to localize objects with varying sizes and shapes (see evidence in Figure 1 (Left) and Figure 6 (a)). Because objects are not always perfectly bounded by fixed-size square patches.
We argue that the inability of standard ViT-UNets to capture spatially heterogeneous representations restricts their scalability and adaptability to various segmentation tasks, especially those involving heterogeneous targets. One notable example is the multi-organ segmentation task, where TransUNet and SwinUNet do not exhibit good scalability when increasing the number of parameters (as evident in Figure 5). In contrast, models that can capture spatially varying representations (i.e., CoTr [41] and MERIT [29]) showed better scalability. However, the building block in the encoder and decoder of CoTr remains convolutional layers with the bottleneck replaced by a Deformable DETER [49]. MERIT [29] only captures multi-scale feature representations by having a multi-resolution input, which may fail to capture feature representation with varying shapes. We kindly direct readers to Appendix A for a more detailed comparison of the proposed method with CoTr and MERIT. Furthermore, the adaptability of ViT-UNet deteriorates when training samples are extremely scarce. This is particularly true in 3D multi-organ segmentation with only 18 training samples, where SwinUNETR [10] and nnFormer [47] did not empirically show superior performance compared to CNN-based nnUNet [17]. We conjecture that the hierarchical feature extraction in CNNs naturally captures structures within objects at multiple scales. However, this is more challenging for fixed-size window attention and patch embedding in SwinUNet and nnFormer.
@@
Contributions: To address limitations in previous ViT-UNet designs, we propose a spatially agile pure ViT-UNet to capture diverse target objects in a medical image segmentation task. We introduce a structured approach to introduce spatially dynamic components that can capture objects with varying sizes and shapes into a standard ViT-UNet. First, we replaced the standard rigid square patch embedding in ViT-UNet with a novel deformable patch embedding. Second, we adopted spatially dynamic self-attention [40] as the building block to capture spatially varying features. Third, we proposed a novel multi-scale deformable positional encoding to model irregularly sampled grids in self-attention. We integrated these dynamic components into a novel ViT-UNet architecture, named AgileFormer.
We validated our method on three medical image segmentation tasks, including multi-organ, cardiac, and brain tumor MRI segmentation. Extensive experiments demonstrated the effectiveness of the proposed method for three medical image segmentation tasks. Specifically, AgileFormer outperformed recent state-of-the-art UNet models for both 2D and 3D medical image segmentation across all three segmentation tasks, demonstrating exceptional scalability. Remarkably, AgileFormer set a new state-of-the-art performance for multi-organ segmentation, achieving an average DSC (%) of 85.74 and 87.43 in 2D and 3D settings with fewer parameters and FLOPs (see Figure 1 (Right)). Importantly, our empirical investigations also revealed critical components (i.e., patch embedding and positional encoding) in designing a ViT-UNet, which was neglected by previous works but has a significant impact on performance.
2 Related work
CNN-based segmentation methods: During the past decade, convolutional neural networks (CNNs) have been considered the defacto standard for medical image segmentation following the emergence of UNet [30]. Many follow-up works extend the U-shaped encoder-decoder network by introducing different skip-connection methods, including attention-based skip-connection (e.g., Attn-UNet [26] and UCTransUNet [36]) and dense skip-connection (e.g., UNet++ [48] and UNet 3+ [13]). Another line of works modified the convolution building block in UNet by introducing residual connection [45, 18], deformable convolution [22], and large kernel convolution [21, 20, 43]. In addition, methods that leverage multi-scale information for capturing representation at varying sizes have been investigated [46, 31]. Previous work also explored a generalized UNet framework, termed nnUNet [17], which provides a self-configuring UNet framework for any dataset. Finally, MedNeXt [32] extended nnUNet by proposing a transformer-like architecture design of nnUNet following the advances in model designs outlined in ConvNeXt [24].
ViT-based segmentation methods: Since its introduction in 2020, ViT [8] has significantly shifted the trajectory of the design of UNets. This is because self-attention in ViTs is more advantageous in capturing long-range dependencies between image patches than convolution in CNNs. In particular, early explorations integrated ViTs into traditional CNN-based UNets by replacing the CNN encoder [5, 11] and bottleneck [38, 41] with a ViT. There are also hybrid models [12, 41, 27, 28], where the main feature extraction in the encoder/decoder is achieved by both CNN and ViT. However, due to the quadratic complexity and the lack of locality in standard self-attention, few pure ViT-based UNets [4, 47, 16, 29] have been proposed for medical image segmentation tasks, with most relying on the shifted window attention [23].
However, prevailing methods focus on advancing the self-attention building block in a ViT-UNet, while neglecting the other important components (e.g., patch embedding and positional encoding) in building a ViT-UNet. In contrast, our empirical findings revealed that the design of the patch embedding and positional encoding were equally important as the self-attention in enhancing the performance of a ViT-UNet. In addition, none of the prior ViT-UNets for medical image segmentation has explored capturing spatially dynamic feature representation to handle segmentation targets of varying sizes and shapes. In contrast, our Agileformer provides a systematical way to capture spatially dynamic feature representation by a novel deformable patch embedding, spatially dynamic self-attention, and deformable positional encoding.
Figure 2: A roadmap going from a SwinUNet to the design of AgileFormer on Synapse dataset. From top to bottom, each row represents a model design variant, including patch embedding, self-attention, and positional encoding. The foreground bars represent DSC (%) in the FLOP (G) regime of different design variants; a hatched bar means the modification results in a performance drop.
Figure 3: The overview of the proposed AgileFormer. AgileFormer is a U-shape vision transformer consisting of deformable patch embedding as well as neighborhood and deformable self-attention building block with a deformable positional encoding. For illustrative purposes, we take the 2D AgileFormer as an example. Please refer to  Section 3.4 for the detailed discussion on extending 2D AgileFormer to 3D AgileFormer for volumetric segmentation tasks.
@@ 3 Method
In this section, we provide a roadmap going from the standard SwinUNet to the proposed AgileFormer (see Figure 2). To make this paper self-contained, we first provide a brief introduction of the essential elements in ViT-UNets. A ViT-UNet (e.g., SwinUNet) is a U-shaped encoder-decoder network with skip connections, wherein the primary feature extractions in both the encoder and decoder are achieved by the self-attention mechanism. The building block of a standard ViT-UNet is composed of three fundamental components: patch embedding, self-attention, and positional encoding. The patch embedding projects image patches into feature embeddings. More recent ViTs [39, 40] even treat downsampling (i.e., patch merging in a SwinUNet) as part of the patch embedding. In this paper, we will follow the same approach. The self-attention, which captures the dependencies between image patches, is used for the main feature extraction. For computational tractability and the need of locality, recent ViT-UNets use window-based self-attention mechanism [4, 12, 15, 10]. Unlike convolution, self-attention discards the spatial correlation between image patches, which hinders the localization ability of pure ViT in segmentation tasks. The positional encoding is used to address this limitation.
@@ 3.1 Deformable Patch Embedding
The ViT-UNet starts by converting image patches into tokens. This process typically involves partitioning an image into a sequence of non-overlapping $n\times n\ (\times n)$ square patches (e.g., $4\times 4$ in SwinUNet [4] for 2D segmentation tasks and $4\times 4\times 4$ in nnFormer for 3D segmentation tasks ). Subsequently, each of these patches is projected into a 1D feature vector. The main reason for performing a rigid (square) patch embedding is due to its simplicity, as it can be easily implemented as a standard convolution (kernel size $k=n$; stride $s=n$; dilation $d=1$). However, we argue that this rigid patch embedding is not optimal for a segmentation task for two main reasons. First, segmentation requires more precise pixel-level localization. The rigid patch embedding can only provide patch-level localization. Second, the shape and size of target objects in most medical image segmentation tasks vary significantly.
Deformable patch embedding. To address the limitation of rigid patch embedding, we propose a deformable patch embedding (see Figure 3) by leveraging the deformable convolution [7]. The deformable convolution is defined as
@@
where $\boldsymbol{f}\in\mathbb{R}^{L\times d_{f}}$ is a $d_{f}$-dimensional feature map with a uniform grid of $L$ locations $p\in\mathbb{R}^{L\times D}$ ($D=2$ for 2D; $D=3$ for 3D). $\boldsymbol{k}$ is the convolutional kernel that operates on the grid $\Omega=[p_{k}]$ and defines the $k$-nearest (i.e., kernel size) neighboring locations of $p$. $[\Delta p]$ are offsets from which the irregular grid is sampled. The offsets are learned through a single convolutional layer $\Delta p=\operatorname{Conv}_{\text{offset}}(\boldsymbol{f})$. $\phi$ is a sampling function that performs bilinear/trilinear interpolation to sample locations $[p+\Delta p+p_{k}]$ in $\boldsymbol{f}$, as the offsets $\Delta p$ are typically fractional.
@@
Downsampling patch embedding. We also replaced the patch merging in standard SwinUNet with convolutional downsampling via a single layer ($k=3$; $s=2$; $d=1$). We used overlapping kernels to better preserve localized patterns [14, 39, 47], which also aligns with the overlapping deformable patch embedding.
@@ 3.2 Spatially Dynamic Self-Attention
The self-attention is the building block of a ViT UNet. Unlike convolution, self-attention does not enforce any spatial inductive bias but makes decisions purely by relying on dependencies (mainly similarities) between tokens while lacking the ability to capture spatially adaptive features for multi-class segmentation. Accordingly, we propose using a spatially dynamic self-attention module as the building block of ViT-UNet. This module is inspired from [39], and it includes deformable multi-head attention (DMSA) [39] and neighborhood multi-head attention (NMSA) [9]. The transformer block was constructed by alternating these two attention mechanisms (see Figure 3). We also distributed more computation to the third encoder block with a stage ratio of $[1:2:5:1]$, instead of $[2:2:2:2]$ as used in [4, 47]. This is because the third layer of the encoder typically captures better feature representations than other layers [36].
3.2.1 Deformable Multi-head Self-Attention
@@
$\operatorname{DMSA}_{h}(\boldsymbol{f})=\operatorname{softmax}(\boldsymbol{Q}_{h}\tilde{\boldsymbol{K}}_{h}^{\top}/\sqrt{d_{k}})\tilde{\boldsymbol{V}}_{h}$ (2)
@@
$\begin{split}\boldsymbol{Q_{h}}=\boldsymbol{f}\boldsymbol{W}_{h}^{Q},\quad\tilde{\boldsymbol{K}}_{h}&=\tilde{\boldsymbol{f}}\boldsymbol{W}_{h}^{K},\quad\tilde{\boldsymbol{V}}_{h}=\tilde{\boldsymbol{f}}\boldsymbol{W}_{h}^{V}\\ \text{with}\quad&\tilde{\boldsymbol{f}}=\phi(\boldsymbol{f};p+\Delta p_{h}).\end{split}$ (3)
@@
Here, we reuse the notation in Equation 1 with $p$ being a uniform grid of points, $\Delta p_{h}$ being the generated offsets for the $h$-th head, and $\phi$ being an interpolation function. $\{\boldsymbol{W}_{h}^{Q},\boldsymbol{W}_{h}^{K}\}\in\mathbb{R}^{d_{f}\times d_{k}}$ and $\boldsymbol{W}_{h}^{V}\in\mathbb{R}^{d_{f}\times d_{v}}$ are trainable parameters, and $d_{k}$, $d_{v}$ are the hidden dimensions of linear projection of the key $\tilde{\boldsymbol{K}}_{h}$ and the value $\tilde{\boldsymbol{V}}_{h}$ in DMSA. The offsets in DMSA are also generated by passing the query through a convolutional layer $\Delta p=\operatorname{Conv}_{\text{offset}}(\boldsymbol{Q}_{h})$. The resulting irregularly sampled feature map is denoted as $\tilde{\boldsymbol{f}}$ in Equation 3. Similar to deformable convolution, the irregularly sampled feature map is then applied to self-attention by using irregularly sampled key $\tilde{\boldsymbol{K}}_{h}$ and value $\tilde{\boldsymbol{V}}_{h}$ (see Equation 2).
3.2.2 Neighborhood Multi-head Self-Attention